# Independence and Induction in Reverse Mathematics

### v2

@@ 7. Nondominated families of functions
$\mathsf{RCA}_{0}\vdash\mathsf{DOM}\rightarrow\mathsf{NDF}$
More interesting is the question whether the converse holds; the following theorem gives a negative answer. This is achieved by separating $\mathsf{DOM}$ from $\mathsf{NDF}$ by showing that, whereas $\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ implies full arithmetical induction as observed in Theorem 6 above, $\mathsf{NDF}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ does not even imply $\mathsf{I}\mathrm{\Sigma}_{3}^{0}$.
@@
$\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}+\mathsf{NDF}$ is $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$.
@@
Repeating the argument, we can obtain $M_{n}$ for all $n<\omega$, and finally we can let $M_{\omega}$ be the model obtained in the limit. Then $\mathcal{M}_{\omega}\models\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}+\mathsf{NDF}$, as desired. ∎
@@ 8. Bi-immunity
Statement 25 (Bi-immunity principle, $\mathsf{BI}$).
@@
The following is a formalization of the well-known fact that every hyperimmune set computes a bi-immune set. Note that, in combination with our previous results, this creates a connection between $\mathsf{BI}$ and $\mathsf{DOM}$, $\mathsf{MAD}$, and induction strength, among others.
@@
$\mathsf{RCA}_{0}+\mathsf{NDF}\vdash\mathsf{BI}$.
@@
Let $\mathcal{M}=(M,\mathcal{S})$ be a model of $\mathsf{NDF}$, and fix any $A\in\mathcal{S}$. It is sufficient to construct a set $B$ such that no $C=\Phi^{A}_{e}$ is contained in $B$ or $\overline{B}$.
Consider the family of use functions of the functionals $\Phi^{A}_{e}$ for every $e$, and let $f$ witness $\mathsf{NDF}$ with respect to that family. In order to decide which elements we want to put into $B$, we proceed as follows for each $y$ in ascending order: Check if there exists an $e<y$ such that $\Phi^{A}_{e,f(y)}(y){\downarrow}$ but such that for all $x<y$, either $\Phi^{A}_{e,f(y)}(x){\uparrow}$ or $\Phi^{A}_{e,f(y)}(x){\downarrow}=B(x)$. If no such $e$’s exist, then let $y\not\in B$; and if they do exist, then pick the least such $e$ and put $y$ into $B$ or $\overline{B}$ in such way as to ensure $B(y)\neq\Phi^{A}_{e}(y)$. By $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$, every total $\Phi_{e}^{A}$ eventually gets treated in this way. ∎
Finally, here is a simple conservation result.
@@
The proof is an imitation of that used by Harrington [6, IX.2] in proving that $\mathsf{WKL}_{0}$ is $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}$.
@@
  • At stages of the form $s+1=2i+2$ for some $i$, let $\sigma_{s+1}$ be the shortest extension of $\sigma_{s}$ which maximizes the size of $\{e<c_{i}\colon\Phi_{e}(e){\downarrow}\}$.

### v3

@@ 7. Hyperimmunity in Reverse Mathematics
$\mathsf{RCA}_{0}\vdash\mathsf{DOM}\rightarrow\mathsf{HI}$. ∎
The reverse direction does not hold, as this example shows.
Example 25.
Consider the $\omega$-model whose second-order part $\mathcal{S}$ consists of all second-order objects whose Turing degree is low for Martin-Löf random. These oracles form a Turing ideal not containing any high Turing degree, [^1] See, for instance, Nies [7, Chapter 5]. thus $\mathsf{DOM}$ does not hold in this $\omega$-model. Furthermore, every set in $\mathcal{S}$ is $\Delta_{2}$ and $\mathcal{S}$ is not topped. Thus for every $A\in\mathcal{S}$ there is a $B\in\mathcal{S}$ with $B\gneq_{T}A$ hyperimmune relative to $A$. Thus $\mathsf{HI}$ holds.
The following theorem confirms the separation of these two notions with an alternative proof; it separates $\mathsf{DOM}$ from $\mathsf{HI}$ by showing that, whereas $\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ implies full arithmetical induction as observed in Theorem 6, $\mathsf{HI}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ does not even imply $\mathsf{I}\mathrm{\Sigma}_{3}^{0}$.
@@
$\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}+\mathsf{HI}$ is $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$.
@@
Repeating the argument, we can obtain $M_{n}$ for all $n<\omega$, and finally we can let $M_{\omega}$ be the model obtained in the limit. Then $\mathcal{M}_{\omega}\models\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}+\mathsf{HI}$, as desired. ∎
@@ 8. Bi-immunity
Statement 27 (Bi-immunity principle, $\mathsf{BI}$).
@@
The following is a formalization of the well-known fact that every hyperimmune set computes a bi-immune set. Note that, in combination with our previous results, this creates a connection between $\mathsf{BI}$ on the one hand and $\mathsf{DOM}$, $\mathsf{MAD}$ and induction strength on the other.
@@
$\mathsf{RCA}_{0}+\mathsf{HI}\vdash\mathsf{BI}$.
@@
Let $\mathcal{M}=(M,\mathcal{S})$ be a model of $\mathsf{HI}$, and fix any $A\in\mathcal{S}$. It is sufficient to construct a set $B$ such that no $C=\Phi^{A}_{e}$ is contained in $B$ or $\overline{B}$.
Consider the family of use functions of the functionals $\Phi^{A}_{e}$ for every $e$, and let $f$ witness $\mathsf{HI}$ with respect to that family. In order to decide which elements we want to put into $B$, we proceed as follows for each $y$ in ascending order: Check if there exists an $e<y$ such that $\Phi^{A}_{e,f(y)}(y){\downarrow}$ but such that for all $x<y$,
$\text{either}\quad\Phi^{A}_{e,f(y)}(x){\uparrow}\quad\text{or}\quad\Phi^{A}_{e,f(y)}(x){\downarrow}=B(x).$
If no such $e$’s exist, then let $y\not\in B$; and if they do exist, then pick the least such $e$ and put $y$ into $B$ or $\overline{B}$ in such way as to ensure $B(y)\neq\Phi^{A}_{e}(y)$. By $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$, every total $\Phi_{e}^{A}$ eventually gets treated in this way. ∎
The converse result does not hold. This can either be shown directly by constructing a model of $\mathsf{RCA}_{0}$ that satisfies $\mathsf{BI}$ but not $\mathsf{HI}$, or by employing the following conservation result.
@@
The proof is an imitation of that used by Harrington [8, IX.2] in proving that $\mathsf{WKL}_{0}$ is $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}$.
@@
  • At stages of the form $s+1=2i+2$ for some $i$, let $\sigma_{s+1}$ be the shortest extension of $\sigma_{s}$ which maximises the size of $\{e<c_{i}\colon\Phi_{e}(e){\downarrow}\}$.